# Edge Weight Prediction For Category-Agnostic Pose Estimation

### v1

@@ Abstract
Category-Agnostic Pose Estimation (CAPE) localizes keypoints across diverse object categories with a single model, using one or a few annotated support images. Recent works have shown that using a pose graph (i.e., treating keypoints as nodes in a graph rather than isolated points) helps handle occlusions and break symmetry. However, these methods assume a static pose graph with equal-weight edges, leading to suboptimal results.
We introduce EdgeCape, a novel framework that overcomes these limitations by predicting the graph’s edge weights which optimizes localization. To further leverage structural priors, we propose integrating Markovian Structural Bias, which modulates the self-attention interaction between nodes based on the number of hops between them. We show that this improves the model’s ability to capture global spatial dependencies. Evaluated on the MP-100 benchmark, which includes 100 categories and over 20K images, EdgeCape achieves state-of-the-art results in the 1-shot setting and leads among similar-sized methods in the 5-shot setting, significantly improving keypoint localization accuracy. Our code is publicly available.
@@ 1 Introduction
Refer to caption Refer to caption
Support Data CapeFormer [36]
(Only Nodes)
Refer to caption Refer to caption
GraphCape [14] Ours
(Unweighted Graph) (Weighted Graph)
Figure 1: EdgeCape. Given a support image, keypoints definition, and skeletal relations (purple), our model localizes the keypoints on a query image (green). Previous methods treat keypoints as isolated (CapeFormer) or use unweighted graphs (GraphCape). We, in contrast, predict weighted graphs that lead to better localization.
@@
2D pose estimation, which involves identifying the locations of key semantic parts within an image, is a fundamental problem in computer vision. From human pose estimation [11, 4, 45] to animal tracking and vehicle localization [37, 30], accurate pose estimation is essential for both academic research and industrial applications. Traditional approaches mainly focused on category-specific models that are tailored to predefined keypoints and predefined object categories. These models achieve high accuracy in some domains but struggle when encountering categories that lack annotated training data. This limitation has sparked interest in flexible models that generalize beyond the fixed categories and keypoints seen in training.
To address these challenges, Category-Agnostic Pose Estimation (CAPE) has emerged as a promising solution [42]. CAPE seeks to enable keypoint localization for arbitrary keypoints and for any object category using only a few annotated support images, allowing a single model to generalize across diverse object types. By significantly reducing the need for extensive data collection and retraining for each new category or keypoint definition, CAPE offers a versatile and cost-effective approach to pose estimation.
Refer to caption Refer to caption
(a) (b)
Figure 2: Graph Definition: (a) Graph structure serves as a strong prior, representing an object’s 3D structure/anatomy, and remains valid across viewing angles for both rigid and non-rigid objects. (b) As Edge placement can be ambiguous, we aim to learn the optimal graph for keypoint localization. We visualize different input graphs (left) and unnormalized output predicted graphs (right). Our network disconnects symmetric parts, which can hurt localization, and converges to similar refined skeletons for both inputs.
While early CAPE methods treat keypoints as isolated entities, recent works [14, 33] have leveraged user-defined structural pose graphs to break symmetry, manage occlusions, and preserve object structure. However, these methods rely on unweighted graphs, limiting their effectiveness. For instance, for locating a human’s right elbow, it is effective to consider the positions of the right hand and shoulder, though their relative contributions may differ in magnitude. These structural priors hold for any viewing angle, and both for rigid and non-rigid objects. We aim to advance this line of work by introducing more complex graphs with real-valued edge weights. However, assigning weights to edges, rather than just defining their presence, is difficult even for humans. In addition, determining the optimal skeletal structure for an object is an ill-posed problem, as different skeletal annotations can lead to varying performance [14]. Our contribution lies in predicting graph edge weights that optimize keypoint localization. Figure 1 highlights the differences between recent CAPE methods.
In this paper, we present EdgeCape, a novel approach to CAPE that extends the graph-based framework by incorporating adaptive edge weight prediction. Predicting the full structure of unseen object categories can be extremely challenging, as it requires a 3D understanding of the object’s structure/anatomy. Instead, we learn to refine input prior graphs and assign edge weights. This approach combines user-provided prior knowledge with predicted edge weight, producing a structure-aware skeleton optimized for localization. This way, our model can handle complex and varied object geometries. During training, we learn edge weights prediction using user-annotated graphs [14]. At test time, a user can provide any prior graph, which we refine to improve keypoint localization. An illustration of different graph priors and predictions is shown in Figure 2.
To further exploit the graph structure for improved localization, we draw on recent advances in Graph Neural Networks [48, 54, 9, 40], which integrate graph structures into the self-attention mechanism of Transformers. Specifically, based on Ma et al. [22] we treat the normalized adjacency matrix as a Markov transition matrix, representing transitions between keypoints as a stochastic process. This probabilistic view allows us to define a bias term that adjusts self-attention weights according to the structural distance between nodes in the graph, supporting more complex spatial relations.
@@
We evaluate our approach on the MP-100 benchmark, a comprehensive dataset comprising over 20,000 images spanning 100 diverse object categories. Our method outperforms the previous state-of-the-art in the 1-shot setting and outperforms similar-sized methods in the 5-shot setting, significantly improving keypoint localization accuracy. We also demonstrate the robustness of our approach under challenging conditions and through extensive ablation study, such as super cross-category keypoint matching, where traditional methods often fail.
@@
In summary, we introduce a novel approach to CAPE that treats object structure as dynamic and adaptive, leveraging predicted edge weights to improve localization accuracy. Our contributions can be summarized as follows:
  • We propose an adaptive edge weight prediction mechanism, enabling the model to learn complex real-valued instance-specific pose graphs, rather than relying on fixed binary skeleton definitions.
@@
  • We propose an enhanced graph-based architecture, with Markovian Attention Bias, allowing the model to better capture complex spatial dependencies between keypoints.
@@ 2.1 Category-Agnostic Pose Estimation
Category-agnostic pose estimation (CAPE), introduced by Xu et al. [42], aims to extend conventional category-specific pose estimation [11, 4, 45, 49, 47] and multi-category pose estimation [43, 49] to unseen categories. This approach focuses on developing models that generalize beyond category-specific training, providing flexibility and robustness across various object types. Several key models have contributed to the field of CAPE. POMNet [42], based on regression-based approaches [18, 25, 56], utilized a transformer to encode query images and support keypoints to predict similarity. Building on this, CapeFormer [35] adopted a DETR-like framework [5, 50, 12, 39] to address unreliable matching outcomes, refining initial predictions.
Later, different approaches were suggested to improve various aspects of CapeFormer. ESCAPE[24], introduced super-keypoints that capture the statistics of semantically related keypoints from different categories, tackling variability in object appearances and poses. Chen et al. [7] proposed to predict meta-points independently of support annotations, later refining them to align with desired keypoints based on support inputs. These meta-points are similar to object proposal tokens [1, 15, 57] in object detection models and provide structural information. X-Pose [44] introduced a bottom-up approach, capable of simultaneously working on several instances in an image. PPM [27] used large-scale text-to-image diffusion models [32] to locate keypoints. At test-time, pseudo-prompts, corresponding to the keypoints, are learned and then used for localization using the diffusion model’s cross-attention maps.
Moreover, recent works found object structure is crucial for effective CAPE localization. GraphCape[14] leveraged graph convolutional networks (GCNs) [16] to model the structural relations between keypoints, to better handle symmetry and occlusions. Later, CapeX [33] expanded this method by using textual point explanations, enabling the model to infer object poses from natural language descriptions. SCAPE [19] suggested enhancing the modeling of structural details by altering keypoints correlations. They refine the attention maps in self-attention blocks by using learnable attention filters.
SDPNet [31] also explored graph-based methods by using GCNs to facilitate information sharing between structural keypoints. Their approach predicts adjacency matrices through keypoints’ self-attention, and is supervised by a secondary GCN along with a mask-reconstruction task for self-supervised learning. However, there are several limitations to this method. First, it relies solely on keypoint features, while object structure might be better inferred from the entire object. Second, predicting structures for unseen categories without prior knowledge is extremely challenging and requires a 3D understanding of the object’s anatomy. Finally, some structural information might be encoded in the auxiliary GCN rather than the adjacency matrix, causing it to be lost as it is only used during training. In contrast, we use both image and keypoint features to refine the user’s prior structural knowledge and leverage the model’s decoder for supervision, constructing a weighted graph optimized for localization.
@@ 2.2 Structural Encoding in GNNs
Graph Transformers adapt the Transformer architecture for graph-structured data. In the context of GNNs, a transformer’s self-attention mechanism can be viewed as a message passing between all nodes, regardless of the graph connectivity. Three main strategies exist for incorporating structural information into Transformers:
Positional Embedding from Graph Structure. Mialon et al. and Feldman et al. [23, 13] introduce positional encodings based on heat kernels and other graph kernels. Dwivedi et al. [9] employed Laplacian eigenvectors as positional encodings. Building upon this, Kreuzer et al. [17] utilized the complete spectrum of the Laplacian matrix to learn more expressive positional encodings. Recently, Ma et al. [22] employed random walk probabilities as relative positional encodings.
Combining Graph Neural Networks (GNNs) with Transformers. Wu et al. [41] used GNNs to extract local structural node features, which a Transformer then processes to model long-range interactions. Rampášek et al. [28] introduced a hybrid model that combines GNN layers with self-attention layers to capture both local and global dependencies. Similarly, Hirschorn et al. [14] replaced the feed-forward layers of the transformers with GCNs.
Incorporating Graph Structural Bias into Self-Attention. This strategy transforms graph structural features into biases that directly modify the self-attention mechanism, allowing the Transformer to capture graph-specific characteristics better. Ying et al. [48] proposed spatial encodings based on the shortest path lengths between nodes to model structural similarity. Zhao et al. [54] enhanced the attention matrix by incorporating proximity-based relationships across different node neighborhoods. Dwivedi et al. [9] extended the graph Transformer model by encoding edge features into the self-attention mechanism. Wu et al. [40] introduced topological information as relational biases, improving the attention matrix’s fidelity.
Among these methods, we advocate incorporating graph structural biases into the self-attention matrix and combining GNNs with Transformers, allowing the Transformer to naturally understand and leverage node connectivity.
Figure 3: Framework Overview. Our model consists of three main components: a feature extraction module, a skeleton predictor, and a graph-based keypoint predictor. The feature extraction module processes multi-scale image features, while the skeleton predictor refines the prior graph input by predicting residual connections. The graph-based keypoint predictor then utilizes the keypoint relations based on the refined structure, improving keypoint localization across diverse object geometries.
@@ 3 Method
The complete framework of our method is illustrated in Figure 3, and is based on GraphCape [14]. In this setup, the model extracts support keypoints and query image features, and refines them by a transformer encoder through self-attention. A similarity proposal generator provides initial coordinate localization, and a graph transformer decoder then refines these initial predictions.
In the following section, we describe the key components of our proposed approach, EdgeCape. We first present the skeleton prediction network, the core idea of our paper. Then, we discuss the supervision strategy for adjacency matrix prediction. Next, we detail our enhanced feature extraction process. Finally, we describe our use of Markovian Attention Bias, which incorporates graph-based biases into the Transformer’s self-attention mechanism. More details on the complete framework and about the training scheme are in the supplementary.
3.1 Edge Weight Prediction
The core idea of our work is to learn adaptive weighted skeletons optimized for localization and tailored to each input instance. A well-designed graph emphasizes localized, relevant connections that adapt to the unique structure of each object. Weighted connections are crucial because they enable the model to capture varying levels of influence between keypoints. For example, complete graphs, where all keypoint are connected, introduce unnecessary complexity and dilute the model’s focus with irrelevant relations. In contrast, structural or anatomical nearest-neighbor connections are beneficial for extrapolating locations, especially in cases of occlusion. For instance, in human pose estimation, a weighted pose graph provides a strong prior for locating a person’s knee based on the thigh and foot positions, though their relative contributions may differ in magnitude. Moreover, training models with structural priors also helps in learning asymmetrical features.
Therefore, instead of relying on unweighted, fixed graphs, we aim to create adaptive weighted skeletons. To achieve this, we predict only the residual graph structure rather than the entire graph. By refining a reliable unweighted graph prior, the model leverages the structural or anatomical relations provided by the user, a feature that becomes especially valuable for unseen object categories with unfamiliar 3D structures.
Formally, we define a learnable function, $f_{\theta}$ to learn the residual graph $\Delta A$. This function leverages input prior skeleton knowledge denoted $A_{prior}$, as well as the support image features $F_{s}$ and support keypoint features $F_{s}^{k}$, both of which are crucial for structure learning. The function $f_{\theta}$ is defined as:

### v2

# EdgeCape: Edge Weight Prediction For Category-Agnostic Pose Estimation

@@ Abstract
Category-Agnostic Pose Estimation (CAPE) localizes keypoints across diverse object categories with a single model, using one or few annotated support images. Recent works have shown that using a pose-graph (i.e., treating keypoints as nodes in a graph rather than isolated points) helps handle occlusions and break symmetry. However, these methods assume a given pose-graph with equal-weight edges, leading to suboptimal results. We introduce EdgeCape, a novel framework that overcomes these limitations by predicting the graph’s edge weights in order to optimize localization. To further leverage structural (i.e., graph) priors, we propose integrating Markov Attention Bias, which modulates the self-attention interaction between nodes based on the number of hops between them. We show that this improves the model’s ability to capture global spatial dependencies. Evaluated on the MP-100 benchmark, which includes 100 categories and over 20K images, EdgeCape achieves state-of-the-art results in the 1-shot and 5-shot settings, significantly improving localization accuracy. Our code will be publicly available.
@@ 1 Introduction
2D pose estimation, which involves identifying the locations of key semantic parts within an image, is a fundamental problem in computer vision. From human pose estimation (Fang et al., 2022a; Cao et al., 2019; Yang et al., 2021) to animal tracking and vehicle localization (Song et al., 2019; Reddy et al., 2018), accurate pose estimation is essential for both academic research and industrial applications. Traditional approaches mainly focused on category-specific models that are tailored to pre-defined keypoints and pre-defined object categories. These models achieve high accuracy in some domains but struggle when encountering categories that lack annotated training data. This limitation has sparked interest in flexible models that generalize beyond the fixed categories and keypoints seen in training.
To address these challenges, Category-Agnostic Pose Estimation (CAPE) has emerged as a promising solution (Xu et al., 2022a). CAPE enables keypoint localization for arbitrary keypoints and any object category using only a few annotated support images, allowing a single model to generalize across diverse object types. By significantly reducing the need for extensive data collection and retraining for each new category or keypoint definition, CAPE provides a versatile and cost-effective approach to pose estimation. However, CAPE remains exceptionally challenging as it must infer keypoint relations that are structurally meaningful across vastly different object categories using a few annotated examples. This is crucial for breaking symmetry, handling occlusions, and preserving object structure. Figure 1 highlights the differences between recent CAPE methods. Although early works treated keypoints as isolated entities, recent works (Hirschorn and Avidan, 2024; Rusanovsky et al., 2024) address this challenge by leveraging user-defined structural graphs. These graphs describe the skeletal relations between keypoints and are termed pose-graphs. These structural priors hold for any viewing angle and both for rigid and non-rigid objects. Graph-based methods provide robustness to occlusions, handle symmetric structures, and enforce anatomical consistency.
While effective, these methods rely on user-provided unweighted pose-graphs, significantly limiting their adaptability. For example, when locating a human’s right elbow, both the right hand and shoulder provide useful context, but their contributions should vary in strength. In conventional pose estimation, these relative contributions are learned by leveraging the fact that the object category is fixed, meaning its pose-graph definition is shared and can be learned from abundant same-category examples. In contrast, CAPE presents fundamentally new and largely unexplored challenges: object categories are unknown, and keypoints are determined dynamically at test time. We aim to advance graph-based CAPE by leveraging optimal pose-graphs with real-valued edge weights. However, multiple challenges arise. First, humans struggle to determine the optimal pose-graph, as demonstrated in Figure 2. This is crucial, as different graph definitions significantly impact localization performance (Hirschorn and Avidan, 2024). Moreover, requiring users to manually specify edge weights is impractical, as there is no clear correct assignment. Second, existing edge prediction methods are unsuitable for CAPE. Self-attention-based approaches, such as Graph Attention Transformers (Veličković et al., 2017), struggle in category-agnostic settings because they fail to model structural keypoint relations (Hirschorn and Avidan, 2024). Alternatively, GCN variants (Cai et al., 2019; Shi et al., 2019) assume a fixed pose-graph - an assumption that does not hold for CAPE. Lastly, CAPE requires generalization to unseen object categories, making direct pose-graph prediction particularly challenging, as it demands an implicit understanding of 3D object structure and anatomy. Thus, a more sophisticated approach is needed.
Refer to caption Refer to caption Refer to caption Refer to caption
Support Data CapeFormer GraphCape EdgeCape (Ours)
(Only Nodes) (Unweighted Graph) (Weighted Graph)
Figure 1: Given a support image, keypoints definition, and skeletal relations (support data) from any category, our model localizes the keypoints on a query image. Previous methods treat keypoints as isolated (CapeFormer (Shi et al., 2023b)) or use unweighted graphs (GraphCape (Hirschorn and Avidan, 2024)). We, in contrast, predict weighted graphs that lead to better localization.
In this paper, we introduce EdgeCape, a novel approach that extends the graph-based CAPE framework by tackling the challenge of category-agnostic pose-graph prediction. Our goal is to find the optimal weighted pose-graph for any class. Rather than inferring the entire pose-graph from scratch, we refine a user-provided unweighted pose-graph by learning to assign edge weights and add or remove connections as needed. This approach eliminates the need for 3D prior knowledge and enables the model to adapt to any object category with varying geometries using a simple structural prior. Furthermore, our method learns subtle instance-specific adjustments, which are crucial to handling CAPE’s inherent ambiguity. In addition, we overcome the lack of ground-truth optimal edge weights through a self-supervised strategy that predicts pose-graphs to maximize localization accuracy and improve occlusion handling. Finally, to better exploit pose-graphs for improved localization, we introduce a Markov attention bias that adjusts the model’s self-attention according to the structural distance between nodes in the graph, supporting more complex spatial relations.
@@
We evaluate our approach on the MP-100 benchmark, a comprehensive dataset comprising over 20,000 images spanning 100 diverse categories. Our method outperforms the previous state-of-the-art, significantly improving localization accuracy. We also demonstrate the robustness of our approach under challenging conditions through an extensive ablation study.
@@
In summary, we introduce a novel CAPE method that enables a more nuanced capture of object geometry, leveraging optimal weighted pose-graphs to improve localization accuracy. Our contributions can be summarized as follows:
  • We are the first to explore category-agnostic pose-graph prediction, introducing a novel mechanism that learns instance-specific weighted graphs that generalize across unseen object categories.
@@
  • We propose an enhanced graph-based CAPE architecture with Markov Attention Bias, allowing the model to better capture complex spatial dependencies between keypoints.
@@
Figure 2: Which of these pose-graphs is optimal? As edge placement can be ambiguous, we aim to learn the optimal pose-graph for category-agnostic keypoint localization. We show several graph annotations for the same image. The second-to-last (cyan) serves as input to our model; the last (orange) is our predicted pose-graph. While not visually superior, it achieves the best performance.
@@ 2.1 Category-Agnostic Pose Estimation
Category-agnostic pose estimation (CAPE), introduced by Xu(Xu et al., 2022a), aims to extend conventional category-specific pose estimation (Fang et al., 2022a; Cao et al., 2019; Yang et al., 2021; Yu et al., 2021; Yang et al., 2022) and multi-category pose estimation (Xu et al., 2022b; Yu et al., 2021) to unseen categories, enabling models that generalize beyond category-specific training for greater flexibility and robustness. POMNet (Xu et al., 2022a), a regression-based approach (Li et al., 2021; Oberweger and Lepetit, 2017; Zimmermann and Brox, 2017), utilized a transformer to encode query images and support keypoints for similarity prediction. Building on this, CapeFormer (Shi et al., 2023a) adopted a DETR-like framework (Carion et al., 2020; Zhang et al., 2022; Fang et al., 2022b; Wang et al., 2022) to address unreliable matching outcomes, refining initial predictions. Later, different approaches were suggested to improve various aspects of CapeFormer. ESCAPE(Nguyen et al., 2024), introduced super-keypoints that capture the statistics of semantically related keypoints from different categories, tackling variability in object appearances and poses. Chen (Chen et al., 2024) predicted meta-points independent of support annotations, later refined to align with target keypoints. Similar to object proposal tokens (Alexe et al., 2012; Hosang et al., 2015; Zitnick and Dollár, 2014), meta-points also provide structural cues. SCAPE (Liang et al., 2024) simplified the task by directly regressing keypoint locations without refinement. X-Pose (Yang et al., 2024) introduced a bottom-up approach for multi-instance localization. PPM (Peng et al., 2024) harnessed Stable Diffusion (Rombach et al., 2022), learning pseudo-prompts for each keypoint at test time and localizing via diffusion model cross-attention maps. This idea was later expanded by Chen et al. (2025b) which explored weak-shot keypoint estimation using intra-image and inter-image correspondences. FMMP (Chen et al., 2025a) employed multiscale features and leveraged denser keypoint localization by randomly mixing structurally linked keypoints. While effective, its reliance on multi-scale feature extraction restricts the range of backbone architectures that can be utilized. CapeLLM Kim et al. (2025) uses text in order to locate keypoints, using LLMs to replace the need for support examples. Recent work highlights the importance of object structure for CAPE localization. GraphCape (Hirschorn and Avidan, 2024) used graph convolutional networks (GCNs) (Kipf and Welling, 2016) to model structural relations between keypoints, improving robustness to symmetry and occlusions. CapeX (Rusanovsky et al., 2024) extended this by integrating pose-graphs with textual point explanations, enabling pose inference from natural language. SDPNet (Ren et al., 2024) also adopted graph-based methods, employing an auxiliary GCN for information sharing between keypoints. It predicts adjacency matrices via keypoint self-attention, supervised by a secondary GCN and a mask-reconstruction task for self-supervised learning. However, these methods face key limitations: (1) they rely only on keypoint features, whereas object structure may be better inferred from the entire object; (2) generalizing to unseen categories without prior knowledge is highly challenging and requires 3D anatomical understanding; and (3) structural cues may be encoded in the auxiliary GCN rather than the adjacency matrix, leading to information loss since the GCN is used only during training. In contrast, we combine image and keypoint features to refine prior structural knowledge and leverage decoder supervision to build a weighted graph optimized for localization. Although category-agnostic graph prediction remains underexplored, AutoLink He et al. (2022) validates the utility of learned weighted graphs for pose estimation. However, it is constrained to a category-specific setting and a fixed skeletal topology. In contrast, CAPE necessitates category-agnostic generalization, adapting to varying object categories and keypoint definitions at test time.
@@ 2.2 Structural Encoding in GNNs
Graph Transformers extend Transformers to graph-structured data, where self-attention can be viewed as message passing between all nodes, independent of connectivity. Structural information is incorporated using three main strategies:
Positional Embedding from Graph Structure. Mialon and Feldman (Mialon et al., 2021; Feldman et al., 2022) introduce positional encodings based on graph kernels. Dwivedi (Dwivedi and Bresson, 2020) used Laplacian eigenvectors, later extended by Kreuzer (Kreuzer et al., 2021) with spectral encodings for greater expressivity. Ma (Ma et al., 2023b) instead modeled adjacency matrices as stochastic processes, using random-walk probabilities as relative encodings.
Integrating Graph Neural Networks (GNNs) with Transformers. Wu (Wu et al., 2021) captured local features with GNNs followed by Transformers to model long-range dependencies. Rampášek (Rampášek et al., 2022) combined GNN and self-attention layers, while Hirschorn (Hirschorn and Avidan, 2024) replaced Transformer feed-forward layers with GCNs.
Incorporating Graph Structural Bias into Self-Attention. Ying (Ying et al., 2021) encodes shortest-path distances as biases that directly alter the attention mechanism. Zhao (Zhao et al., 2021) incorporated proximity-based neighborhood relations. Dwivedi (Dwivedi and Bresson, 2020) integrated edge features into attention, while Wu et al. (Wu et al., 2022) injected topological information as relational biases to enhance attention fidelity.
Figure 3: Framework Overview. Our model consists of two main components: a pose-graph predictor (visualized in Figure 15) and a graph-based keypoint predictor. The pose-graph predictor refines the prior graph input by predicting residual connections. The graph-based keypoint predictor then utilizes the predicted keypoint relations, improving localization across diverse object structures.
@@ 3 Method
The full framework of our method is shown in Figure 3. In the following, we outline the key components of EdgeCape: (1) our graph-based CAPE formulation, (2) the category-agnostic pose-graph prediction network, and (3) an enhanced graph decoder that embeds structural biases. Additional details are in the supplementary (Section C).
3.1 Preliminaries
CapeFormer (Shi et al., 2023b) treats keypoints as individual entities, sharing information via self-attention. While effective for single-category pose estimation, it struggles to generalize structural understanding to unseen categories (Hirschorn and Avidan, 2024). Consequently, self-attention alone is insufficient to capture structural relations for CAPE. GraphCape (Hirschorn and Avidan, 2024) addresses this limitation by incorporating user-provided pose-graphs as a structural prior. The pose-graph is represented by an adjacency matrix $A_{prior}\in\{0,1\}^{K\times K}$, where $a_{ij}=1$ if node $v_{i}$ is connected to node $v_{j}$, and 0 otherwise. Using graph convolutions, information is propagated between connected keypoints, aiding in structure preservation and occlusion handling. Additionally, integrating graphs during training helps break feature symmetry by enforcing similarity based on structural relations rather than purely semantic attributes. Our work is built on GraphCape, and we include the full framework details in the supplementary (Section C.1).
3.2 Category-Agnostic Pose-Graph Prediction
Category-agnostic pose-graph prediction presents a fundamentally new and largely unexplored challenge: object structures vary widely across categories, making it difficult to infer structure from scratch. Our key insight is to predict adjustments to a user-provided prior $A_{\text{prior}}$ rather than learning the entire graph. This yields expressive, weighted pose-graphs that adapt to object-specific structures while remaining grounded in prior knowledge. Formally, we define a learnable function $f_{\theta}$ to learn the residual pose-graph in the form of an adjacency matrix $\Delta A\in\mathbb{R}^{K\times K}$: